# SWAN: SGD with Normalization and Whitening Enables Stateless LLM Training

### v2

@@ 4.2 Practical Considerations
Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse. We hence directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation. This is given by (Song et al., 2022; Li et al., 2018):
${\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}),\quad{\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}$ (3)
where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$. At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2). See Appendix H for implementation details.
@@
Note that the same N-S procedure has been discussed in (Mei et al., 2023) and  (Jackson, 2023) for preconditioned optimizers. Another N-S procedure has also been introduced in (Bernstein & Newhouse, 2024b), and optimized in (Jordan et al., 2024) under BF16 precision.
Computational Overhead.
[…]
Rescaling $\mathtt{GradWhitening}$.
@@
The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, and as such might drastically change the effective learning rate. In practice, we propose the following re-scaling before updating the weight:
$\Delta{\mathbf{W}}^{(t)}\leftarrow\frac{\|\tilde{{\mathbf{G}}}^{(t)}\|\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}$ (4)
This helps rescale the norm of the whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates; we found that the learning rates used by Adam tend to work well for SWAN. We can remove rescaling, but doing so would require additional hyper-parameter search to find the optimal learning rate.
@@ 5.2 $\mathtt{GradNorm}$: stabilizing gradient distributions of STB
Consider the STB (Definition 1). Assuming we inherit the assumptions in Theorem 1 of Tian et al. (2023), as described in Appendix B. Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole. Then, its standardized stochastic gradients $\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}({\bm{x}}^{(t)})}{{\mathbf{U}}_{C}^{\mathbf{W}}})$ satisfy:
@@ 6 Experiments
In this section, we report empirical results for SWAN. The main purposes are:
  • Benchmarking the performance, training speed-ups and memory footprint of SWAN on pre-training LLaMA with C4 dataset (Raffel et al., 2020) against Adam. Results are presented in Section 6.1, Section 6.2, and Section 6.3.
  • Investigating the contribution of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ to the performance of SWAN through ablation studies, as outlined in Section 6.4.
  • Separately evaluating whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ achieve their intended purposes of stabilizing gradient noise and modeling/ counteracting local curvature, respectively. Detailed results are reported in Section 6.5.1 and Section 6.5.2.
All experiments run on NVIDIA A100 GPUs.
6.1 SWAN Performance on LLM Pre-training Tasks
@@
We evaluate SWAN on large language model (LLM) pre-training tasks using a LLaMa-based architecture Touvron et al. (2023a) with RMSNorm and SwiGLU activations (Zhang & Sennrich, 2019; Gao et al., 2023). We consider models with 60M, 130M, 350M, and 1.3B parameters, all trained on the C4 dataset Raffel et al. (2020) using an effective token batch size of 130K tokens (batch size 512, context length 256). Following the setup of Zhao et al. (2024a), SWAN is applied to all linear modules in both attention and MLP blocks. Training uses BF16 precision by default. The other evaluation settings are the same as Zhao et al. (2024a).
Table 1: Comparison with Adam and its memory-efficient low-rank variants on pre-training various sizes of LLaMA models on C4 dataset. Validation perplexity is reported, along with a memory estimate of the total of parameters and optimizer states based on BF16 format. Since we used train scripts and configs based on the code opensourced by Zhao et al. (2024a), all baseline results are taken from Zhao et al. (2024a). The exceptions are the results for Adam and Galore, as we cannot exactly reproduce the results reported in Zhao et al. (2024a); hence we reported results from our own runs instead.
@@
SWAN 32.28 (0.12G) 24.13 (0.25G) 18.22 (0.68G) 15.13 (2.60G)
Adam 33.02 (0.36G) 24.44 (0.76G) 20.11 (2.06G) 16.44 (7.80G)
Galore 33.09 (0.24G) 24.67 (0.52G) 19.74 (1.22G) 15.89 (4.38G)
Low-Rank 78.18 (0.26G) 45.51 (0.54G) 37.41 (1.08G) 142.53 (3.57G)
@@
SWAN speed up vs Adam 1.2 X 1.23 X > 2X > 2X
@@
Figure 4: Comparison of convergence rate of different methods on LLM pretraining tasks.
@@
Since SWAN aims to explore EMA-free adaptive optimizers for LLMs, we focus on comparing it with Adam and other low-rank optimizers. The baselines are:
  • Adam (Kingma & Ba, 2015): one standard choice for training large models.
  • Galore Zhao et al. (2024a): a memory-efficient Adam variant with low-rank gradient projections.
  • Low-Rank (Kamalakara et al., 2022): a low-rank factorization approach (${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$).
  • LoRA Hu et al. (2021): the LoRA method applied for pre-training as in Zhao et al. (2024a).
  • ReLoRA Lialin et al. (2023): a full-rank extension of LoRA with parameter merging.
Hyperparameters
We use the publicly available hyperparameter configurations from Zhao et al. (2024a) for all baselines [^2] This does not include Adam, since the authors of Zhao et al. (2024a) did not disclose their Adam configs. However, we used the same learning rate sweep procedure as described in their paper. This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models., including method-specific learning rates selected through a grid search. For SWAN, we use a default setting without extensive tuning. We match the global learning rate (0.001) used by Adam, except for the 1.3B model (Adam diverges at 0.001, see Appendix H). We use 10 steps of Newton–Schulz iteration for $\mathtt{GradWhitening}$. Unlike Adam, SWAN does not require a warm-up phase to ensure stable training. All other scheduling parameters remain the same as in the baselines. Full details can be found in Appendix H.
@@
We follow Kaddour et al. (2024) by allocating the same total training budget (in terms of processed tokens) for all methods. By the end of training, all methods decay to the same target learning rate. We also leave SWAN’s learning rates largely untuned for a more direct comparison.
@@ Results
Table 1 and Figure 4 show the results. SWAN achieves lower validation perplexities than the baselines with lower estimated memory consumptions. On the 350M and 1.3B models, SWAN reaches over a 2$\times$ speedup (in steps or tokens) relative to Adam (Figure 1 and Figure 5 (a)). Further details are in Section 6.2.
6.2 Is the Improvement Multiplicative or Additive?
A key question in assessing speedup factors is whether the improvement over Adam is multiplicative or additive. A multiplicative speedup implies that the optimizer’s relative advantage remains proportionally consistent over time, while an additive speedup suggests a less desired constant step advantage. We explore these scenarios using two plots, the speed-up ratio comparison and the perplexity comparison (Figure 5), across different model sizes.
@@ 6.2 Is the Improvement Multiplicative or Additive?
where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ are the training steps required by Adam and SWAN, respectively, to reach perplexity $P$.
@@
This represents the expected perplexity of SWAN if it only consistently outperforms Adam by $\Delta$ steps.

### v3

@@ 4.2 Practical Considerations: Fast and Stable Implementation
Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse. One option is to directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation. This is given by (Song et al., 2022; Li et al., 2018):
$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k})\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}\end{cases}$ (3)
@@
where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$. At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2). See Appendix J for implementation details. Note that the same N-S procedure has been discussed in (Mei et al., 2023) and  (Jackson, 2023) for preconditioned optimizers. Another N-S procedure was been introduced in (Bernstein & Newhouse, 2024b), and optimized in (Jordan et al., 2024).
However, estimating $({\mathbf{G}}{\mathbf{G}}^{\top})^{-1/2}$ with NS requires $\mathcal{O}(m^{3})$ (assuming $m<n$) complexity, making its scalability for larger models, especially under distributed settings unclear Jordan et al. (2024). Here, we propose a heuristic scheme that has $\mathcal{O}(m^{2})$ complexity to estimate square-root inverse (comparable to the $\mathcal{O}(mn)$ complexity of Adam second moment):
$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}\text{Diag}(3{\mathbf{I}}-{\mathbf{Z}}_{k}\text{Diag}({\mathbf{Y}}_{k}))\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-\text{Diag}({\mathbf{Z}}_{k}){\mathbf{Y}}_{k})\text{Diag}({\mathbf{Z}}_{k})\end{cases}$ (4)
where under a slight abuse of notation, $\text{Diag}(\cdot)$ returns a diagonal matrix that has the same diagonals as the input matrix. Intuitively, whenever we encounter matrix multiplication in NS iterations, we replace one of the matrices with its diagonal approximation. The particular choice of where to apply $\text{Diag}(\cdot)$ is determined by performance on synthetic datasets. We refer to this as the NS with diagonal substitution (NSDS) scheme. In Appendix B, we empirically demonstrate that NSDS performs well in minimizing the matrix condition number of gradients. As shown in Figure 7, on the gradient distribution induced by Llama model training, the performance of NSDS (when combined with $\mathtt{GradNorm}$) is comparable to or better than the standard NS scheme in reducing the condition number.
We found that NSDS does have the side effect of slowing down the early stage convergence. However, it offers stronger tail convergence in return. For LLM pretraining NSDS helps SWAN achieve similar or better final validation loss performance than the original NS scheme (Section 6.1), enabling Adam-level training throughput without the help of distributed computation (Section 6.2).
Rescaling for robustness.
@@
The operator $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, and as such might drastically change its effective learning rate. In practice, we propose the following re-scaling before updating the weights:
$\Delta{\mathbf{W}}^{(t)}\leftarrow\frac{\|\tilde{{\mathbf{G}}}^{(t)}\|\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}=\frac{\sqrt{mn}\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}$ (5)
This helps rescale the norm of the whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, that is $\|\tilde{{\mathbf{G}}}^{(t)}\|=\sqrt{mn}$. Notice that this is exactly the norm of signed gradient descent ($=$ Adam without EMAs) (Bernstein et al., 2018). This relationship allows us to directly adapt the learning rates from Adam without tuning specifically for SWAN. This can be throught as a variation of LR grafting (Agarwal et al., ).
Finally, the practical SWAN algorithm ($\mathtt{GradNorm}$ + NSDS-$\mathtt{GradWhitening}$ + rescaling) is presented in Algorithm 1 (under $\mathtt{diag}=\text{True}$). We denote it as SWAN.
4.3 Saving analysis: balancing memory efficiency and expressiveness
Table 1 shows the theoretical memory savings of SWAN. When applied on a 2D tensor of $m$ by $n$ ($m>n$), SWAN requires storing 00 optimizer states (i.e., intermediate variables maintained over time), compared with other state-of-the-art memory efficient optimizers. Notably, this is achieved without significantly compromising its capabilities to model the adaptive learning rates. SWAN can model element-wise adaptive learning rates, while the concurrent work (SGD-sal and Apollo) needs to resort to block-wise or even tensor-wise approximations. In our experiments in Section 6.1, we show that these approximations indeed has an negative impact on LLM training performance.
@@ 5.2 $\mathtt{GradNorm}$: stabilizing gradient distributions of STB
Consider the STB (Definition 1). Assuming we inherit the assumptions in Theorem 1 of Tian et al. (2023), as described in Appendix C. Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole. Then, its standardized stochastic gradients $\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}({\bm{x}}^{(t)})}{{\mathbf{U}}_{C}^{\mathbf{W}}})$ satisfy:
@@ 5.3 $\mathtt{GradWhitening}$: An Efficient Non-diagonal Second-Order Update
5.4 Verifying the Theoretical Insights
We empirically validate the theoretical benefits of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. As detailed in Appendix I, we demonstrate that $\mathtt{GradNorm}$ effectively stabilizes stochastic gradient distributions during SGD training of a scaled-down LLaMA model on the C4 dataset, evidenced by significantly reduced KL divergence fluctuations compared to standard training. Additionally, $\mathtt{GradWhitening}$ significantly enhances optimization performance across high-dimensional and ill-conditioned classic optimization functions, consistently outperforming both SGD and Adam. These findings confirm that $\mathtt{GradNorm}$ promotes gradient stability and $\mathtt{GradWhitening}$ addresses local curvature challenges, enabling faster and more reliable performance.
@@ 6 Experiments
In this section, we report empirical results for SWAN. All experiments run on NVIDIA A100 GPUs.
6.1 Memory-Efficient Pre-training of Llama Models
@@
We evaluate SWAN on memory-efficient Llama  (Touvron et al., 2023a) pre-training tasks, using the standardized settings of (Zhao et al., 2024a), which has been adopted by many recent works (Zhao et al., 2024a; Chen et al., 2024; Zhang et al., 2024c; Huang et al., ; Zhu et al., 2024). We consider models with 60M, 130M, 350M, and 1.3B parameters, all trained on the C4 dataset Raffel et al. (2020) using an effective batch size of 130K tokens. Following the setup of Zhao et al. (2024a), SWAN is applied to all linear modules in both attention and MLP blocks. Training uses BF16 by default unless specified, see Appendix J. The other evaluation settings follows Zhao et al. (2024a).
Table 2: Comparison with Adam and its memory-efficient low-rank variants on pre-training various sizes of LLaMA models on C4 dataset. Validation perplexity is reported, along with a memory estimate of the total of parameters and optimizer states based on BF16 format. Baseline results are directly taken from the official numbers reported in Zhao et al. (2024a); Zhu et al. (2024), as they shares exactly the same setup. Note that for Adam we report both the official results from Zhao et al. (2024a) and our reproduced result.
@@
Adam 33.02 (0.32G) 24.44 (0.75G) 19.24 (2.05G) 16.44 (7.48G)
Adam (cited) 34.06 (0.32G) 25.08 (0.75G) 18.80 (2.05G) 15.56 (7.48G)
SWAN 30.59 (0.23G) 22.61 (0.43G) 16.63 (0.93G) 13.56 (2.98G)
SWAN 30.00 (0.23G) 22.83 (0.43G) 17.14 (0.93G) 14.42 (2.98G)
SWAN-0 32.28 (0.23G) 24.13 (0.43G) 18.22 (0.93G) 15.13 (2.98G)
Momentum$+$$\mathtt{GradWhitening}$ 31.6 (0.27G) 24.59 (0.59G) 19.30 (1.49G) 16.08 (5.23G)
Galore 34.88 (0.26G) 25.36 (0.57G) 18.95 (1.29G) 15.64 (4.43G)
Apollo-mini 31.93 (0.23G) 23.53 (0.43G) 17.18 (0.93G) 14.17 (2.98G)
Apollo 31.55 (0.26G) 22.94 (0.57G) 16.85 (1.29G) 14.20 (4.43G)
Low-Rank 78.18 (0.26G) 45.51 (0.57G) 37.41 (1.29G) 142.53 (4.43G)
@@
SWAN speed up vs Adam 1.52 X 1.6 X > 2.3 X > 2.4 X
@@
Figure 4: Comparison of convergence rate of different methods on LLM pretraining tasks. The training curves of Adam, Galore and Apollo-mini are reproduced according to the opensource code of Zhao et al. (2024a); Zhu et al. (2024). We further compare to their official results in Table 2.
SWAN optimizer
We consider three variants: SWAN-0, which aims to show-case the robustness and effectiveness of our method out-of-the-box, with almost no tuning. It uses naive NS-iteration for whitening, disabled learning rate warmup, and use similar learning rates optimized for Adam. SWAN, the vanilla version of our method, in which we enabled learning rate warmup, and allowed the use of optimized learning rates that largely differ from Adam. Finally, SWAN, the strongest and most efficient variant that employs the proposed NSDS scheme for fast whitening (section 4.2). For all SWAN variants, we adopt a lazy-tuning approach (hyperparameters are set without extensive search) to reduce the possibility of unfair performance distortion. Notably, for SWAN, we share the same hyperparameters across all model sizes. Full details can be found in Appendix J.
@@
We focus on comparing SWAN with Adam (Kingma & Ba, 2015) and other memory-efficient optimizers. All baseline results are directly quoted from corresponding papers as they all share the same standard setup. The baselines include Adam (Kingma & Ba, 2015), which is a standard choice for training large models; and Galore Zhao et al. (2024a), a memory-efficient Adam variant with low-rank gradient projections. We also consider Low-Rank (Kamalakara et al., 2022), a low-rank factorization approach (${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$); and LoRA Hu et al. (2021), which applies the LoRA method for pre-training as in Zhao et al. (2024a). Additionally, we include ReLoRA Lialin et al. (2023), a full-rank extension of LoRA with parameter merging, and Momentum$+$$\mathtt{GradWhitening}$, which applies Newton-Schulz $\mathtt{GradWhitening}$ on top of momentum instead of $\mathtt{GradNorm}$; this is equivalent to Muon Jordan et al. (2024) with Nesterov acceleration turned off. Finally, we compare with Apollo-mini and Apollo (Zhu et al., 2024). Full details can be found in Appendix J.
@@
We follow Kaddour et al. (2024) by allocating the same total training budget (in terms of tokens) for all methods. By the end of training, all methods decay to the same target learning rate, which is $10\%$ of the initial learning rates. This approach eliminates an often over-looked evaluation bias, where optimizers with earlier learning rate cool-downs tend to have an unfair advantage due to artificially fast convergence caused by learning rate decay Kaddour et al. (2024).
@@ Results
As shown in Table 2 and Figure 1, all SWAN variants achieve strong performance compared with baselines, requiring the lowest memory consumptions comparable to SGD. Across all models, SWAN-0 surpasses the performance of the Adam and Momentum-$\mathtt{GradWhitening}$ (muon-like) baseline in terms of validation perplexities. SWAN further delivers a comparable performance to Apollo series; and finally, SWAN with NSDS scheme delivers the strongest, SOTA-level performance under this setup. Notably, on the 350M and 1.3B models, all SWAN variants reaches at least 2$\times$ speedup (in steps or tokens) relative to Adam (Figure 1 and Figure 5 (a)). Finally, we observe that in general, SWAN has a slower convergence speed in the early stage than SWAN. However, it offers strong tail convergence in return and surpasses SWAN in terms of final performance.
6.2 Memory Efficiency and Throughput Analysis
Table 3: Raw and effective throughput analysis, under different model parallelization (MP) settings.
Method Raw / eff. throughput (1B)
Adam 117872 / 117872 (tokens/s)
SWAN w/o MP   58600 / 117200 (tokens/s)
SWAN w/o MP 107808 / 258739 (tokens/s)
SWAN w/ MP 114160 / 228320 (tokens/s)
SWAN w/ MP 115872 / 278092 (tokens/s)
Memory Footprint
We compare SWAN, Adam, and Galore on a single A100 GPU. Unlike Zhao et al. (2024a); Zhu et al. (2024), which report layer-wise training memory usage, we measure total end-to-end memory consumption under full-model training using a batch-size of $1$ for 1.3B, 7B, and 13B models. As shown in Figure 1 (c), SWAN’s memory usage is on par with SGD, providing nearly a 50% reduction in total memory. If we further incorporate the per-layer training technique of (Zhao et al., 2024a; Lv et al., 2023), SWAN further achieves $\approx 70\%$ total memory reduction on top of $\approx 100\%$ reduction on optimizer states. This underlines the benefit of the stateless design.
Effective Throughput
We assess throughput when training a 1.3B LLama model on 8 A100 GPUs with a batch size of 130K. All gradient processing are done in BF16. We use two metrics: raw throughput: number of tokens processed per second. Effective throughput: raw throughput adjusted by SWAN’s token efficiency relative to Adam. These metrics evaluate the impact of different $\mathtt{GradWhitening}$ schemes on training speed, and also account for the fact that some optimizers make more effective use of training tokens. As shown in Table 3, SWAN with naive NS $\mathtt{GradWhitening}$ requires model parallelization (where NS of different tensors are distributed to different GPUs) to achieve competitive throughput. Without model parallelization, the throughput of SWAN is 50 % lower than Adam. With the proposed NSDS scheme, SWAN achieves an raw throughput comparable to Adam, without any need for model parallelization. Consequently, SWAN exhibits a 2 $\times$ higher effective throughput than Adam.
6.3 Is the Speed-up Multiplicative or Additive?
A key question regarding speedup factors is whether the improvement over Adam is multiplicative or additive. A multiplicative speedup implies that the optimizer’s relative advantage remains proportionally consistent over time, while an additive speedup suggests a less desired constant step advantage. To investigate this, we take the SWAN-0 setup as an example, and address this question using two plots, the speed-up ratio comparison and the perplexity comparison (Figure 5), across model sizes.
@@ 6.3 Is the Speed-up Multiplicative or Additive?
where $S_{\text{Adam}}(P)$/$S_{\text{SWAN{}}}(P)$ are the training steps required by Adam/SWAN to reach perplexity $P$.
@@
This represents the expected perplexity of SWAN if it only outpaces Adam by fixed $\Delta$ steps.